# From Generation to Detection: A Multimodal Multi-Task Dataset for Benchmarking Health Misinformation

### v1

@@ 4 Experiments
GPT4o GPT4o-mini Llama3.2-V LLaVA-1.6 Qwen2-VL GPT4o GPT4o-mini Llama3.2-V LLaVA-1.5 Qwen2-VL
ZS ZS ZS FS ZS FS ZS FS ZS ZS ZS FS ZS FS ZS FS
Text Human 0.499 0.499 0.500 0.499 0.500 0.494 0.499 0.499 0.334 0.328 0.246 0.376 0.077 0.348 0.162 0.348
AI 0.499 0.500 1.000 0.436 1.000 0.498 0.500 0.499 0.183 0.105 0.020 0.238 0.010 0.101 0.019 0.113
Text & Image Human 0.499 0.499 0.499 0.499 0.500 0.495 0.499 0.498 0.371 0.353 0.312 0.373 0.085 0.351 0.140 0.340
AI 0.498 0.499 0.498 0.454 1.000 0.498 0.499 0.500 0.314 0.261 0.250 0.331 0.030 0.207 0.084 0.247
AI-T 25% 0.499 0.499 0.499 0.468 1.000 0.497 0.498 0.500 0.347 0.304 0.287 0.349 0.065 0.267 0.124 0.302
AI-T 50% 0.500 1.000 0.499 0.477 1.000 0.497 0.499 0.500 0.347 0.304 0.286 0.347 0.065 0.267 0.124 0.302
AI-T 75% 1.000 1.000 1.000 0.490 1.000 0.497 0.499 0.500 0.395 0.372 0.330 0.381 0.086 0.339 0.207 0.392
AI-I 25% 0.498 0.499 0.498 0.450 1.000 0.499 0.500 0.500 0.313 0.262 0.245 0.328 0.040 0.221 0.089 0.244
AI-I 50% 0.499 1.000 0.499 0.450 1.000 0.499 0.500 0.499 0.310 0.264 0.253 0.333 0.043 0.219 0.075 0.231
AI-I 75% 1.000 1.000 0.499 0.449 1.000 0.497 1.000 0.499 0.310 0.265 0.252 0.333 0.037 0.212 0.077 0.244
@@ 4.2 Baseline Models
We established various multimodal baselines, including both proprietary and open-source models for a comprehensive comparison. Proprietary models: We employed GPT-4o and GPT-4o Mini, which have demonstrated excellent performance on open-source benchmarks such as MMLU Hendrycks et al. (2021), HumanEval Chen et al. (2021a), and MMMU Yue et al. (2024), and are widely adopted in vision and language research. These models serve as a benchmark standard, representing the top zero-shot performance that current state-of-the-art VLLMs can achieve. Open-sourced models: We include three open-source VLLMs for a comprehensive comparison. These models are: Llama-3.2-Vision, built on the Llama-3.1-8B text model with a 2-billion-parameter vision encoder; LLaVA-1.6, which utilises the 7-billion-parameter Qwen v1.5 model with CLIP-ViT-L-336 attached via an MLP projection layer; and Qwen2-VL, which employs Multimodal Rotary Position Embedding (M-RoPE) with a 675M vision encoder and Qwen2-7B, resulting in a total of 7.6B parameters. These models demonstrate competitive zero-shot performance on vision-and-language benchmarks. See appendix for detailed experiment setup and prompts.
@@ 5 Result
GPT4o GPT4o-mini Llama3.2-V LLaVA-1.6 Qwen2-VL GPT4o GPT4o-mini Llama3.2-V LLaVA-1.6 Qwen2-VL
ZS ZS ZS FS ZS FS ZS FS ZS ZS ZS FS ZS FS ZS FS
Text Human 0.316 0.323 0.023 0.321 0.242 0.321 0.493 0.324 0.254 0.266 0.012 0.255 0.201 0.267 0.320 0.288
AI 0.093 0.074 0.327 0.057 0.211 0.041 0.003 0.046 0.113 0.122 0.495 0.083 0.220 0.062 0.003 0.043
Text & Image Human 0.234 0.304 0.140 0.312 0.137 0.319 0.321 0.317 0.263 0.239 0.069 0.237 0.138 0.279 0.301 0.298
AI 0.155 0.119 0.188 0.122 0.271 0.070 0.093 0.071 0.215 0.233 0.231 0.164 0.282 0.063 0.091 0.099
AI-T 25% 0.234 0.172 0.261 0.144 0.261 0.091 0.122 0.094 0.285 0.335 0.259 0.145 0.272 0.079 0.122 0.089
AI-T 50% 0.268 0.188 0.316 0.173 0.239 0.118 0.143 0.125 0.306 0.339 0.284 0.159 0.270 0.091 0.111 0.097
AI-T 75% 0.307 0.205 0.338 0.171 0.203 0.127 0.162 0.138 0.389 0.400 0.305 0.221 0.173 0.096 0.134 0.116
AI-I 25% 0.236 0.143 0.256 0.173 0.262 0.100 0.114 0.096 0.261 0.264 0.279 0.176 0.295 0.067 0.093 0.118
AI-I 50% 0.270 0.180 0.284 0.215 0.235 0.109 0.143 0.108 0.369 0.282 0.319 0.253 0.193 0.082 0.117 0.188
AI-I 75% 0.322 0.218 0.333 0.246 0.185 0.134 0.156 0.130 0.358 0.275 0.316 0.203 0.154 0.093 0.118 0.194
@@
Task 1: Information reliability check. The results of the information reliability check are presented in Table 3. We divide the dataset into reliable and unreliable articles and apply VLLMs for binary classification across various data settings, including text-only, text-image, and mixed text-image content incorporating both human and AI-generated inputs. To assess the impact of AI-generated content, we introduce varying proportions of AI-generated context at 25%, 50%, and 75%, in both text and image modalities. Baseline models perform better at classifying reliable articles than unreliable ones, with a significant disparity in F1 scores observed in both zero-shot and five-shot settings. In zero-shot setting, some models, such as Llava-1.6, GPT-4o, and GPT-4o-min, achieve perfect F1 scores for reliable articles but struggle to classify unreliable articles. This may be caused by that the models tend to classify input information as reliable without proper verification. In the five-shot setting, models’ performance improve on unreliable articles but worsen on reliable ones. This suggests that example-based learning helps the models better understand content for verification.
Task 2: Information originality check. The results of the originality check are presented in Table 4. We classify the dataset into three categories: human-generated, AI-generated, and a mix of text-image content combining human and AI inputs, for multi-label classification. VLLMs are tasked with distinguishing whether the input context is human, AI, or human-AI mixed. Similar to Task 1, we introduce AI-generated content in both text and images at proportions of 25%, 50%, and 75% to further evaluate the robustness of VLLMs in distinguishing mixed content. Baseline models struggle to accurately classify the originality of articles in both zero-shot and five-shot settings. We observe that baseline models may perform worse in the five-shot setting, suggesting that current VLLMs may lack the ability to learn comprehensive representations for assessing information originality using examples. GPT-4o performs the best among all the models, with Llama-3.2-Vision ranking as the second-best model, slightly trailing GPT-4o in performance. This performance difference is likely attributed to the larger dataset and model size used in GPT-4o’s training.
@@ 6 Conclusion
We introduce MM-Health, a multimodal dataset designed for detecting human and AI generated health misinformation. Unlike datasets limited to human generated information or augmented with LLM generated text, our approach combines text and image generative models to create multimodal counterparts of misinformation. To ensure diversity in generative model outputs, we incorporate five different open-source models for each modality. Our experiments involve five different VLLMs, including proprietary GPT-4o and GPT-4omini models, as well as SOTA open-source models, to evaluate data reliability and originality classification. Experimental results reveal that current VLLMs cannot accurately detect misinformation and AI-generated content, highlighting challenges in the era of generative AI. We hope publicly releasing the MM-Health dataset guides future research.
@@
In this work, we limited data collection to two modalities due to their prevalence on mainstream social platforms. This decision restricts the range of misinformation addressed by MM-Health, particularly from short-video platforms like TikTok, thereby limiting the scope of its potential applications. We encourage future studies to incorporate a broader variety of modalities to further enhance the diversity of machine-generated health information. Additionally, we exclusively employed out-of-the-box multimodal language models for misinformation detection, aiming to emulate common practices among the general public as our baseline. Given the existence of several specialised models designed for misinformation detection, we suggest further research to facilitate a more comprehensive comparison.
@@ Appendix A Remove Images
Appendix B Overall Statics
@@ Appendix D Generation Prompts
Appendix E Experiment Settings and prompts
Experiment Settings. For the GPT-4o and GPT-4o mini models, we use the OpenAI API endpoints: gpt-4o-2024-08-06 and gpt-4o-mini-2024-07-18, respectively, throughout our experiments. Open-source models are hosted on a server equipped with an Nvidia A6000 GPU with 48GB of VRAM. To replicate realistic scenarios in most cases, we retain the default parameter settings for all models. To fully utilise the potential of the models, we experiment with baseline models using both zero-shot and five-shot settings, employing consistent prompts designed for different tasks. Zero-shot is directly evaluated on the testing set, while five-shot uses five samples from the training set and evaluates the models on the testing set. We use standard macro F1 scores to evaluate the baseline models across all three tasks. Since our evaluation tasks involve three different data settings with human and AI context, we employ both binary classification (for human and AI) and multi-label classification (for human, AI, and mixed content).
Experiment prompts. We also present the prompts used for all three tasks in the experiment, covering both zero-shot and five-shot settings. Both proprietary and open-source models follow the same prompts. Figure 8 illustrates the zero-shot and five-shot prompts for the information reliability check, Figure 9 depicts the prompts for the information originality check, and Figure 10 presents the prompts for the fine-grained AI detection analysis. Article inputs are highlighted in green, while model outputs are highlighted in orange. Since models may not always produce the exact output specified in the prompts, we run each model several times until all responses are correctly formatted to meet the output requirements.

### v2

@@ 4 Experiments
4o 4o-mini Llama3.2-V LLaVA-1.6 Qwen2-VL MedGemma 4o 4o-mini Llama3.2-V LLaVA-1.5 Qwen2-VL MedGemma
ZS ZS ZS FS ZS FS ZS FS ZS FS ZS ZS ZS FS ZS FS ZS FS ZS FS
Text Human 0.499 0.499 0.500 0.499 0.500 0.494 0.499 0.499 0.499 0.485 0.334 0.328 0.246 0.376 0.077 0.348 0.162 0.348 0.295 0.409
AI 0.499 0.500 1.000 0.436 1.000 0.498 0.500 0.499 1.000 0.499 0.183 0.105 0.020 0.238 0.010 0.101 0.019 0.113 0.081 0.208
Text &
Image
Human 0.499 0.499 0.499 0.499 0.500 0.495 0.499 0.498 0.527 0.495 0.371 0.353 0.312 0.373 0.085 0.351 0.140 0.340 0.304 0.393
AI 0.498 0.499 0.498 0.454 1.000 0.498 0.499 0.500 0.532 0.499 0.314 0.261 0.250 0.331 0.030 0.207 0.084 0.247 0.229 0.327
AI-T 25% 0.499 0.499 0.499 0.468 1.000 0.497 0.498 0.500 0.499 0.499 0.347 0.304 0.287 0.349 0.065 0.267 0.124 0.302 0.274 0.381
AI-T 50% 0.500 1.000 0.499 0.477 1.000 0.497 0.499 0.500 0.499 0.499 0.347 0.304 0.286 0.347 0.065 0.267 0.124 0.302 0.274 0.385
AI-T 75% 1.000 1.000 1.000 0.490 1.000 0.497 0.499 0.500 0.499 0.499 0.395 0.372 0.330 0.381 0.086 0.339 0.207 0.392 0.363 0.437
AI-I 25% 0.498 0.499 0.498 0.450 1.000 0.499 0.500 0.500 0.499 0.499 0.313 0.262 0.245 0.328 0.040 0.221 0.089 0.244 0.231 0.333
AI-I 50% 0.499 1.000 0.499 0.450 1.000 0.499 0.500 0.499 0.500 0.500 0.310 0.264 0.253 0.333 0.043 0.219 0.075 0.231 0.261 0.341
AI-I 75% 1.000 1.000 0.499 0.449 1.000 0.497 1.000 0.499 1.000 1.000 0.310 0.265 0.252 0.333 0.037 0.212 0.077 0.244 0.253 0.336
@@ 4.2 Baseline Models
We established various multimodal baselines, including both proprietary and open-source models for a comprehensive comparison. Proprietary models: We employed GPT-4o and GPT-4o Mini, which have demonstrated excellent performance on open-source benchmarks such as MMLU mmmlu, HumanEval chen2021codex, and MMMU MMMU, and are widely adopted in vision and language research. These models serve as a benchmark standard, representing the top zero-shot performance that current state-of-the-art VLLMs can achieve. Open-sourced models: We include three general open-source VLLMs for a comprehensive comparison. These models are: Llama-3.2-Vision, built on the Llama-3.1-8B text model with a 2-billion-parameter vision encoder; LLaVA-1.6, which utilises the 7-billion-parameter Qwen v1.5 model with CLIP-ViT-L-336 attached via an MLP projection layer; and Qwen2-VL, which employs Multimodal Rotary Position Embedding (M-RoPE) with a 675M vision encoder and Qwen2-7B, resulting in a total of 7.6B parameters. To further access the performance of domain specific model, we also include MedGemma, which is a 4B model of Gemma3 variants that are trained for performance on medical text and image comprehension. These models demonstrate competitive zero-shot performance on vision-and-language benchmarks.
4.3 Experiment Settings.
For the GPT-4o and GPT-4o mini models, we use the OpenAI API endpoints: gpt-4o-2024-08-06 and gpt-4o-mini-2024-07-18, respectively, throughout our experiments. Open-source models are hosted on a server equipped with an Nvidia A6000 GPU with 48GB of VRAM. To replicate realistic scenarios in most cases, we retain the default parameter settings for all models. To fully utilise the potential of the models, we experiment with baseline models using both zero-shot and five-shot settings, employing consistent prompts designed for different tasks. Zero-shot is directly evaluated on the testing set, while five-shot uses five samples from the training set and evaluates the models on the testing set. We use standard macro F1 scores to evaluate the baseline models across all three tasks. See appendix for details.
@@ 5 Result
4o 4o-mini Llama3.2-V LLaVA-1.6 Qwen2-VL MedGemma 4o 4o-mini Llama3.2-V LLaVA-1.6 Qwen2-VL MedGemma
ZS ZS ZS FS ZS FS ZS FS ZS FS ZS ZS ZS FS ZS FS ZS FS ZS FS
Text Human 0.316 0.323 0.023 0.321 0.242 0.321 0.493 0.324 0.001 0.075 0.254 0.266 0.012 0.255 0.201 0.267 0.320 0.288 0.006 0.023
AI 0.093 0.074 0.327 0.057 0.211 0.041 0.003 0.046 0.000 0.166 0.113 0.122 0.495 0.083 0.220 0.062 0.003 0.043 0.028 0.203
Text &
Image
Human 0.234 0.304 0.140 0.312 0.137 0.319 0.321 0.317 0.008 0.012 0.263 0.239 0.069 0.237 0.138 0.279 0.301 0.298 0.009 0.014
AI 0.155 0.119 0.188 0.122 0.271 0.070 0.093 0.071 0.156 0.174 0.215 0.233 0.231 0.164 0.282 0.063 0.091 0.099 0.211 0.178
AI-T 25% 0.234 0.172 0.261 0.144 0.261 0.091 0.122 0.094 0.199 0.215 0.285 0.335 0.259 0.145 0.272 0.079 0.122 0.089 0.279 0.241
AI-T 50% 0.268 0.188 0.316 0.173 0.239 0.118 0.143 0.125 0.253 0.283 0.306 0.339 0.284 0.159 0.270 0.091 0.111 0.097 0.308 0.267
AI-T 75% 0.307 0.205 0.338 0.171 0.203 0.127 0.162 0.138 0.280 0.309 0.389 0.400 0.305 0.221 0.173 0.096 0.134 0.116 0.272 0.262
AI-I 25% 0.236 0.143 0.256 0.173 0.262 0.100 0.114 0.096 0.207 0.218 0.261 0.264 0.279 0.176 0.295 0.067 0.093 0.118 0.262 0.235
AI-I 50% 0.270 0.180 0.284 0.215 0.235 0.109 0.143 0.108 0.246 0.256 0.369 0.282 0.319 0.253 0.193 0.082 0.117 0.188 0.259 0.276
AI-I 75% 0.322 0.218 0.333 0.246 0.185 0.134 0.156 0.130 0.291 0.283 0.358 0.275 0.316 0.203 0.154 0.093 0.118 0.194 0.266 0.294
@@
Task 1: Information Reliability Check. The results of the information reliability check are presented in Table 3. We divide the dataset into reliable and unreliable articles and apply VLLMs for binary classification across various data settings, including text-only, text-image, and mixed text-image content incorporating both human and AI-generated inputs. To assess the impact of AI-generated content, we introduce varying proportions of AI-generated context at 25%, 50%, and 75%, in both text and image modalities. Baseline models perform better at classifying reliable articles than unreliable ones, with a significant disparity in F1 scores observed in both zero-shot and five-shot settings. In zero-shot setting, some models, such as Llava-1.6, MedGemma, GPT-4o, and GPT-4o-min, achieve perfect F1 scores for reliable articles but struggle to classify unreliable articles. This may be caused by that the models tend to classify input information as reliable without proper verification. In the five-shot setting, models’ performance improve on unreliable articles but worsen on reliable ones. This suggests that example-based learning helps the models better understand content for verification. MedGemma’s noticeably higher zero-shot and five-shot performance than other open-source models suggests that a model with health-domain-specific knowledge is better equipped to identify the reliability of health-related information.
Task 2: Information Originality Check. The results of the originality check are presented in Table 4. We classify the dataset into three categories: human-generated, AI-generated, and a mix of text-image content combining human and AI inputs, for multi-label classification. VLLMs are tasked with distinguishing whether the input context is human, AI, or human-AI mixed. Similar to Task 1, we introduce AI-generated content in both text and images at proportions of 25%, 50%, and 75% to further evaluate the robustness of VLLMs in distinguishing mixed content. Baseline models struggle to accurately classify the originality of articles in both zero-shot and five-shot settings. We observe that baseline models may perform worse in the five-shot setting, suggesting that current VLLMs may lack the ability to learn comprehensive representations for assessing information originality using examples. GPT-4o performs the best among all the models, with MedGemma ranking as the second-best model, slightly trailing GPT-4o in performance. This performance difference is likely attributed to the larger dataset and model size used in GPT-4o’s training as well as health specific fine-tuning in MedGemma training.
@@
6 Case Study
6.1 Task 1: Information reliability check
Models demonstrated a tendency to rely on superficial stylistic cues rather than semantic content, leading to predictable errors in both directions.
Unreliable Content Misclassified as Reliable. Models were consistently deceived by unreliable articles that mimicked credible journalism. These articles successfully evaded detection by anchoring misleading narratives to a verifiable fact, adopting a professional tone, and avoiding obvious red flags. Instead of containing clear falsehoods, they relied on biased framing and strategic omissions, which the models failed to identify.
Reliable Content Misclassified as Unreliable. Conversely, factually accurate content was often misclassified as unreliable, primarily due to a lack of explicit citations. Models also applied overly cautious heuristics to sensitive domains like medicine and politics, incorrectly penalizing content for using generalized language or offering reasonable predictive analysis.
6.2 Task 2: Information originality check
Similar patterns were observed in detecting text provenance, where models struggled with content that either perfectly emulated a known format or deviated from a perceived natural baseline.
AI-Generated Content Misclassified as Human. AI-generated text evaded detection by adhering closely to formal structures, such as news reports or scientific summaries. Its plausibility was enhanced by incorporating real-world details, like quoting public figures or referencing known events, causing its impersonal and structured format to be mistaken for authentic human writing.
Human-Written Content Misclassified as AI. Simultaneously, human-authored content was frequently misclassified as AI-generated. This was particularly common with text that was concise, logically structured, or satirical. The models appeared to equate clarity and logical structure with machine-generated text, and they failed to grasp rhetorical nuances like irony, often processing satirical arguments literally.
@@ 7 Conclusion
We introduce MM-Health, a multimodal dataset designed for detecting human and AI generated health misinformation. Unlike datasets limited to human generated information or augmented with LLM generated text, our approach combines text and image generative models to create multimodal counterparts of misinformation. To ensure diversity in generative model outputs, we incorporate five different open-source models for each modality. Our experiments involve six different VLLMs, including proprietary GPT-4o and GPT-4omini models, as well as SOTA open-source models, to evaluate data reliability and originality classification. Experimental results reveal that current VLLMs cannot accurately detect misinformation and AI-generated content, highlighting challenges in the era of generative AI. For future work, we plan to explore more robust detection techniques and expand the dataset to include other modalities such as video and audio. We hope publicly releasing the MM-Health dataset guides future research.
Acknowledgements
This research was supported by the Macquarie University Research Acceleration Scheme (MQRAS) and Data Horizon funding.
@@ Acknowledgements
In this work, we limited data collection to two modalities due to their prevalence on mainstream social platforms. This decision restricts the range of misinformation addressed by MM-Health, particularly from short-video platforms like TikTok, thereby limiting the scope of its potential applications. We encourage future studies to incorporate a broader variety of modalities to further enhance the diversity of machine-generated health information. Additionally, we exclusively employed out-of-the-box multimodal language models for misinformation detection, aiming to emulate common practices among the general public as our baseline. Given the existence of several specialised models designed for misinformation detection, we suggest further research to facilitate a more comprehensive comparison. Finally, there is a potential risk that the synthetic health misinformation could be misused by malicious actors. While we acknowledge this risk, the relatively small scale of the data generated in this study is unlikely to significantly exacerbate the existing challenge posed by the vast amount of health misinformation already circulating on the internet.
@@ Appendix A Remove Images
Appendix B Overall Statistics
@@ Appendix D Generation Prompts
Appendix E Experiment Prompts
We present the prompts used for all three tasks in the experiment, covering both zero-shot and five-shot settings. Both proprietary and open-source models follow the same prompts. Figure 8 illustrates the zero-shot and five-shot prompts for the information reliability check, Figure 9 depicts the prompts for the information originality check, and Figure 10 presents the prompts for the fine-grained AI detection analysis. Article inputs are highlighted in green, while model outputs are highlighted in orange. Since models may not always produce the exact output specified in the prompts, we run each model several times until all responses are correctly formatted to meet the output requirements.